### v1

# On the Identification of Latent Objectives in Stochastic Control

Abstract
We propose a variational formulation of an inverse problem in continuous-time stochastic control, aimed at identifying control costs consistent with a given distribution over trajectories. The formulation is based on minimizing the suboptimality gap of observed behavior. We establish a connection between the inverse problem and a generalized dynamic Schrödinger problem, showing that their optimal values coincide. This result links inverse stochastic control with stochastic optimal transport, offering a new conceptual viewpoint on inverse inference in controlled diffusions.
Key words: Inverse problems, stochastic control
I Introduction
Inverse stochastic control concerns the identification of latent control objectives that rationalize observed behavior (see, e.g., Kalman [10], Bellman and Kalaba [2], Ng and Russell [19], Dvijotham and Todorov [7], and Nakano [17]; see also Ab Azar et.al [1], a recent review paper). In continuous-time settings, this problem naturally arises when one seeks to explain a given distribution over state–control trajectories by an underlying optimal control formulation. Despite its relevance to data-driven control and inverse reinforcement learning, a general framework for such inverse problems remains limited.
In this paper, we propose a variational formulation of inverse stochastic control in continuous time based on the suboptimality gap of observed behavior. Given a prescribed distribution over trajectories and a class of candidate running and terminal costs, we quantify how far the observed behavior deviates from optimality under each candidate objective, and define the inverse problem as minimizing this gap. This formulation allows the analysis of observed distributions without assuming optimality of the underlying control policy.
Consider the following forward stochastic control problem:
(I.1) $\inf_{u\in\mathcal{U}}J(u;f,g)$
where
$J(u;f,g)=\mathbb{E}\left[\int_{0}^{T}f(t,X_{t}^{u},u_{t})dt+g(X_{T}^{u})\right].$
The state process is described by the controlled diffusion
(I.2) $dX_{t}=b(t,X_{t},u_{t})dt+\sigma(t,X_{t},u_{t})dB_{t}.$
The control processes are assumed to take values in a closed set $U$ of $\mathbb{R}^{d_{1}}$. The precise definition of the class $\mathcal{U}$ of control processes is given in the Section II below.
Suppose that the probability law $\mu_{t}$ of some state-control process $(X_{t}^{*},u_{t}^{*})$ is given for each time $t\in[0,T]$. Then the problem is to find suitable objective functions $f$ and $g$ such that $\{u_{t}^{*}\}$ is optimal for (I.1).
Let $\mathcal{D}$ be a given set of $(f,g)$ of functions such that $f$ is Borel measurable function on $0,T]\times\mathbb{R}^{d}\times U$ and $g$ is Borel measurable function on $\mathbb{R}^{d}$. For $(f,g)\in\mathcal{D}$, the functions $f$ and $g$ are interpreted as latent running cost and terminal cost functions in the forward optimal control problem (I.1), respectively. Then, to analyze the inverse problem, we introduce the functional
$V(f,g):$ $=\int_{0}^{T}\int_{\mathbb{R}^{d}\times U}f(t,x,u)\mu_{t}(dxdu)dt$
$\quad+\int_{\mathbb{R}^{d}}g(x)\tilde{\mu}_{T}(dx)-J^{*}(f,g)$
where $\tilde{\mu}_{t}$ denotes the probability law of $X_{t}^{*}$ and
(I.3) $J^{*}(f,g)=\inf_{u\in\mathcal{U}}J(u;f,g).$
If $X_{t}^{*}$ is governed by the controlled SDE above, i.e., if $X_{t}^{*}=X_{t}^{u^{*}}$ then
$V(f,g)\geq 0$
for any $(f,g)\in\mathcal{D}$. Moreover, if $u^{*}$ is optimal to the forward control problem (I.1) for some $(f^{*},g^{*})\in\mathcal{D}$, then
$V(f^{*},g^{*})=\inf_{(f,g)\in\mathcal{D}}V(f,g)=0.$
This means that the minimization problem
(I.4) $V^{*}:=\inf_{(f,g)\in\mathcal{D}}V(f,g)$
can be a measure of the identification. From an intuitive viewpoint, $V(f,g)$ quantifies the sub-optimality gap of the observed behavior under the candidate cost $(f,g)$. If the gap vanishes, the observed trajectory is indeed optimal for $(f,g)$. Hence, minimizing $V(f,g)$ provides a natural criterion for model identification, analogous to minimizing a loss function in statistical learning.
The main contribution of the paper is a novel connection between this inverse control formulation and a generalized dynamic Schrödinger problem. For the Schrödinger problem we refer to, e.g, Schrödinger [21] and [22], Bernstein [3], Jamison [8, 9], Dai Pra [5], Mikami [12], Nagasawa [16], Nelson [18], Zambrini [24], and the works cited therein. We show that the optimal value of the inverse problem coincides with that of an stochastic control problem with relaxed marginal constraints. This result establishes a variational link between inverse stochastic control and stochastic optimal transport, providing a new conceptual perspective on inverse inference in controlled diffusions. We also establish existence of optimal solutions to the inverse problem under mild assumptions and discuss additional structural characterizations.
The remainder of the paper is organized as follows. Section II introduces the problem formulation and establishes existence results. Section III presents the connection with generalized Schrödinger-type problems and derives the main variational equivalence.
II Existence
First, we collect some notation used in this paper. Denote by $\mathcal{P}(\mathcal{X})$ the set of all Borel probability measures $\nu$ on a Polish space $\mathcal{X}$. We write $a^{\mathsf{T}}$ for the transpose of a vector or matrix $a$. For Euclidean spaces $\mathcal{X}$ and $\mathcal{Y}$, denote by $C(\mathcal{X},\mathcal{Y})$ the space of all continuous functions $\varphi:\mathcal{X}\to\mathcal{Y}$. In particular, we denote $C(\mathcal{X})=C(\mathcal{X},\mathbb{R})$. We also denote by $C_{b}(\mathcal{X})$ the space of bounded continuous functions on $\mathcal{X}$. Given a Polish space $\mathcal{X}$, for any $\pi\in\mathcal{P}(\mathcal{X})$ and any Borel measurable function $\varphi:\mathcal{X}\to\mathbb{R}$ that is integrable with respect to $\pi$, we put
$\langle\varphi,\pi\rangle:=\int_{\mathcal{X}}\varphi(x)\pi(dx).$
Let $\{B_{t}\}_{0\leq t\leq T}$ be an $m$-dimensional standard Brownian motion on a complete probability space $(\Omega,\mathcal{F},\mathbb{P})$. Further, let $X_{0}$ be a random variable that is independent of $\{B_{t}\}_{0\leq t\leq T}$ such that
$\mathbb{E}|X_{0}|^{2}<\infty.$
Denote by $\{\mathcal{F}_{t}\}_{0\leq t\leq T}$ the augmented filtration generated by $\{B_{t}\}_{0\leq t\leq T}$ and $X_{0}$. For simplicity, we assume $U\ni 0$. The class $\mathcal{U}$ of controls is then defined by the set of all $U$-valued $\{\mathcal{F}_{t}\}$-adapted processes $\{u_{t}\}_{0\leq t\leq T}$ satisfying
$\mathbb{E}\int_{0}^{T}|u_{t}|^{2}dt<\infty.$
To discuss the uniqueness, we impose the following:
  1. (A1)
    There exists a constant $C_{0}>0$ and a modulus of continuity $\rho:[0,\infty)\to[0,\infty)$ such that for $\varphi(t,x,u)=b(t,x,u)$, $\sigma(t,x,u)$ we have
    $|\varphi(t,x,u)-\varphi(t,x^{\prime},u^{\prime})|$ $\leq C_{0}|x-x^{\prime}|+\rho(|u-u^{\prime}|),$
    $|\varphi(t,0,0)|$ $\leq C_{0},$
    for $t\in[0,T]$, $x,x^{\prime}\in\mathbb{R}^{d}$, and $u,u^{\prime}\in U$.
Under (A1), for any given $u\in\mathcal{U}$, there exists a unique solution $\{X_{t}^{u}\}_{0\leq t\leq T}$ of the controlled stochastic differential equation (I.2) with initial condition $X_{0}$ such that
(II.1) $\mathbb{E}\left[\sup_{0\leq t\leq T}|X_{t}^{u}|^{2}\right]<\infty.$
Denote by $P_{t}^{u}$ the law of $(X_{t}^{u},u_{t})$ for each $t\in[0,T]$, i.e., $P_{t}^{u}=\mathbb{P}(X_{t}^{u},u_{t})^{-1}$. Also denote by $\tilde{P}_{t}^{u}$ the law of $X_{t}^{u}$, i.e., $\tilde{P}_{t}^{u}(A)=P_{t}^{u}(A\times U)$, $A\in\mathcal{B}(\mathbb{R}^{d})$. With these notation,
$V(f,g)=\int_{0}^{T}\langle f(t,\cdot,\cdot),\mu_{t}\rangle dt+\langle g,\tilde{\mu}_{T}\rangle-J^{*}(f,g).$
Proposition II.1.
Let $(A1)$ hold. Suppose that $\mu_{t}$ is the marginal law of $(X_{t}^{u^{*}},u^{*}_{t})$ for each $t\in[0,T]$, where $u^{*}=\{u^{*}_{t}\}_{0\leq t\leq T}\in\mathcal{U}$ is an optimal policy to the problem (I.1) for some $(f^{*},g^{*})\in\mathcal{D}$. Then, this $(f^{*},g^{*})$ is a minimizer for the inverse problem (I.4) and we have
$V(f^{*},g^{*})=\inf_{(f,g)\in\mathcal{D}}V(f,g)=0.$
In the following analysis, we do not assume optimality of $\mu_{t}$. Even more, each $\mu_{t}$ may not be necessarily a marginal measure of $(X_{t}^{u},u_{t})$ for some $u\in\mathcal{U}$. However, we do assume the following conditions on $\mu_{t}$ and $\mathcal{D}$:
  1. (A2)
    The family $\{\mu_{t}\}_{0\leq t\leq T}$ satisfies
    $\int_{0}^{T}\int_{\mathbb{R}^{d}\times U}(|x|^{2}+|u|^{2})\mu_{t}(dxdu)dt<\infty.$
  1. (A3)
    There exists a constant $K>0$ such that for any $(f,g)\in\mathcal{D}$ we have
    $|f(t,x,u)|+|g(x)|\leq K(1+|x|^{2}+|u|^{2}),$
    for $t\in[0,T]$, $x\in\mathbb{R}^{d}$, $u\in U$.
Under (A1)–(A3), by (II.1) it is straightforward to check that $V^{*}=\inf_{(f,g)\in\mathcal{D}}V(f,g)>$-$\infty$.
Further, we assume the following:
  1. (A4)
    $U$ is compact.
For $(f,g)\in\mathcal{D}$, define $\phi:[0,T]\times\mathbb{R}^{d}\times U\to\mathbb{R}^{2}$ by
$\phi_{f,g}(t,x,u)=(f(t,x,u),g(x))^{\mathsf{T}}.$
Then we assume the following:
  1. (A5)
    The class $\{\phi_{f,g};(f,g)\in\mathcal{D}\}$ is contained in $C([0,T]\times\mathbb{R}^{d}\times U,\mathbb{R}^{2})$ and is equi-uniformly continuous.
Example II.2.
Assume that each $h=f,g$ is given by
$h(\xi)=\sum_{j=1}^{N}\gamma_{j}e^{-\alpha_{j}|\xi-\xi_{j}|^{2}},\quad\xi\in[0,T]\times\mathbb{R}^{d}\times U,$
for some $N\in\mathbb{N}$, $\gamma_{j}\in\mathbb{R}$, $\alpha_{j}>0$, $\xi_{j}\in[0,T]\times\mathbb{R}^{d}\times U$, $j=1,\ldots,N$, such that $\sum_{j=1}^{N}\left\{|\gamma_{j}|+\alpha_{j}(1+|\xi_{j}|)\right\}$ is bounded by a given positive constant. Then the class $\mathcal{D}$ satisfies (A3) and (A5).
Now we show the existence of a solution to the proposed inverse problem.
Proposition II.3.
Suppose that (A1)(A1)$(A5)$ hold. Then there exists a minimizer of the inverse problem (I.4).
Proof.
Let $\{(f_{n},g_{n})\}_{n=1}^{\infty}\subset\mathcal{D}$ be a minimizing sequence, i.e., $\lim_{n\to\infty}V(f_{n},g_{n})=V^{*}$. By (A3) and (A5), the class $\{\phi_{f,g}:(f,g)\in\mathcal{D}\}$ is compact in $C([0,T]\times\mathbb{R}^{d}\times U,\mathbb{R}^{2})$. It follows from the Ascoli–Arzelà theorem that there exist $(f^{*},g^{*})\in\mathcal{D}$ and $\{n_{k}\}$ such that $\sup_{(t,x,u)\in[0,T]\times B_{R}\times U}|f_{n_{k}}(t,x,u)-f^{*}(t,x,u)|+\sup_{x\in B_{R}}|g_{n_{k}}(x)-g^{*}(x)|\to 0$, $k\to\infty$, for any $R>0$, where $B_{R}$ stands for the closed ball in $\mathbb{R}^{d}$ centered at 0 with radius $R$. Then, the dominated convergence theorem yields $\int_{0}^{T}\langle f_{n_{k}}(t,\cdot,\cdot),\mu_{t}\rangle\,dt\to\int_{0}^{T}\langle f^{*}(t,\cdot,\cdot),\mu_{t}\rangle\,dt$ and $\langle g_{n_{k}},\tilde{\mu}_{T}\rangle\to\langle g^{*},\tilde{\mu}_{T}\rangle$ as $k\to\infty$. Further, applying Theorem 12 in Section 1 of Chapter 3 in Krylov [11], we get
$\varepsilon_{k}$ $:=\sup_{u\in\mathcal{U}}\mathbb{E}\int_{0}^{T}|f_{n_{k}}(t,X_{t}^{u},u_{t})-f^{*}(t,X_{t},u_{t})|\,dt$
$\qquad+\mathbb{E}|g_{n_{k}}(X_{T}^{u})-g^{*}(X_{T}^{u})|\to 0,$
as $k\to\infty$. Therefore, $\lim_{k\to\infty}V(f_{n_{k}},g_{n_{k}})=V(f^{*},g^{*})$. Thus $(f^{*},g^{*})\in\mathcal{D}$ is optimal. ∎
Example II.4.
We consider a one-dimensional stochastic control system governed by the equation
$dX^{u}_{t}=u_{t}\,dt+\frac{1}{10}\,dB_{t},$
where $X_{0}$ is drawn from the standard normal distribution, and $B$ is a standard Brownian motion. The objective is to minimize the cost functional
$J(u;f,g)=\mathbb{E}\int_{0}^{1}\left(10|X^{u}_{t}|^{2}+\theta|u_{t}|^{2}\right)dt,$
where the admissible control space is $U=\mathbb{R}$, and $\theta>0$ is a parameter. This is a special instance of the linear quadratic control framework, for which the value function $v$ admits a closed-form expression. To evaluate the effectiveness of our method, we simulate $10^{4}$ sample trajectories $(X^{(j)}_{t_{i}},u^{(j)}_{t_{i}})$ for $j=1,\ldots,10000$ and $t_{i}=i/1000$ with $i=0,\ldots,1000$, using the above analytical formula for the optimal control with the true parameter $\theta=1$. These synthetic samples are treated as observed data. This example is taken from [17], where the true parameter is estimated by solving a root-finding problem. Here, instead, we frame the estimation of $\theta$ as a stochastic optimization task. We minimize the empirical approximation of the cost functional,
$\frac{1}{10000}\sum_{j=1}^{10000}\sum_{i=0}^{999}\left(10|X^{(j)}_{t_{i}}|^{2}+\theta|u^{(j)}_{t_{i}}|^{2}\right)\Delta t-v(0,X^{(j)}_{0};\theta),$
over $\theta\in[0.0001,10]$, using the Adam optimizer with mini-batch updates. Over 100 independent runs, the average of the estimated values of $\theta$ is found to be $0.9976$, with a standard deviation of $6.0583\times 10^{-3}$.
We remark that some uniqueness results are already obtained in [17]. A regularization of our inverse problems is a subject of a future study.
III Connection with stochastic optimal transport
In this section, we give a connection between (I.4) and stochastic optimal transport such as the Schrödinger problem.
We shall impose the following condition on $\mathcal{D}$:
  1. (B1)
    The class $\mathcal{D}$ is represented as
    $\mathcal{D}=(f_{0},g_{0})+\bigcup_{\lambda>0}\lambda\tilde{\mathcal{D}}$
    where $f_{0}\in C_{b}([0,T]\times\mathbb{R}^{d}\times U)$, $g_{0}\in C_{b}(\mathbb{R}^{d})$, and the class $\{\phi_{f,g};(f,g)\in\tilde{\mathcal{D}}\}$ is non-empty, convex, symmetric at origin, and compact in $C_{b}([0,T]\times\mathbb{R}^{d}\times U,\mathbb{R}^{2})$.
Example III.1.
A class of neural networks satisfies (B1).
For any $\mathcal{P}(\mathbb{R}^{d}\times U)$-valued processes $\pi=\{\pi_{t}\}_{0\leq t\leq T}$ and $\nu=\{\nu_{t}\}_{0\leq t\leq T}$, put
$\rho(\pi,\nu)$ $=\sup_{(f,g)\in\tilde{\mathcal{D}}}\bigg\{\int_{0}^{T}\langle f(t,\cdot,\cdot),\pi_{t}-\nu_{t}\rangle\,dt$
$\quad+\langle g,\tilde{\pi}_{T}-\tilde{\nu}_{T}\rangle\bigg\}.$
Here, $\tilde{\pi}_{t}(dx)=\pi_{t}(dx\times U)$ and $\tilde{\nu}_{t}$ is similarly defined. Note that under (B1), $\rho$ is a pseudometric on the set of $\mathcal{P}(\mathbb{R}^{d}\times U)$-valued processes. Consider the family of approximate bridge problems
$\inf_{u\in\mathcal{U}_{\rho,\varepsilon}}J(u;f_{0},g_{0}),\quad\varepsilon\geq 0,$
where $\mathcal{U}_{\rho,\varepsilon}=\{u\in\mathcal{U}:\rho(P^{u},\mu)\leq\varepsilon\}$. Unlike the classical Schrödinger bridge, the present formulation allows relaxed marginal constraints measured by the pseudometric $\rho$, which is induced by the class of admissible latent objectives.
Then we have the following:
Theorem III.2.
Suppose that $(A1)$ and $(B1)$ hold. Suppose moreover that
(III.1) $\inf_{u\in\mathcal{U}_{\rho,0}}J(u;f_{0},g_{0})<\infty$
and that for any $(f,g)\in\mathcal{D}$ there exists a minimizer $u_{f,g}\in\mathcal{U}$ for the forward control problem (I.1). Then, $V^{*}$ is finite and
$V^{*}$ $=\int_{0}^{T}\langle f_{0}(t,\cdot,\cdot),\mu_{t}\rangle\,dt+\langle g_{0},\tilde{\mu}_{T}\rangle$
$\quad-\lim_{\varepsilon\searrow 0}\inf_{u\in\mathcal{U}_{\rho,\varepsilon}}J(u;f_{0},g_{0}).$
Proof.
Step (i). Take an arbitrary element $(f_{1},g_{1})\in\tilde{\mathcal{D}}$ and $\lambda>0$. Since $f_{0}$, $f_{1}$, $g_{0}$, and $g_{1}$ are all bounded, we have $J^{*}(f_{0}+\lambda f_{1},g_{0}+\lambda g_{1})>$-$\infty$ and so $V^{*}<\infty$.
Let $u_{0}\in\mathcal{U}$ be such that $\rho(P^{u},\mu)=0$. Then, since $\tilde{D}$ is symmetric at origin, for any $(f,g)\in\tilde{\mathcal{D}}$,
$\int_{0}^{T}\langle f(t,\cdot,\cdot),\mu_{t}-P_{t}^{u_{0}}\rangle\,dt+\langle g,\tilde{\mu}_{T}-\tilde{P}_{T}^{u_{0}}\rangle=0.$
Thus, for any $(f,g)\in\tilde{\mathcal{D}}$,
$\int_{0}^{T}\langle f_{0}+\lambda f,\mu_{t}\rangle\,dt+\langle g_{0}+\lambda g,\tilde{\mu}_{T}\rangle$
$-J^{*}(f_{0}+\lambda f,g_{0}+\lambda g)$
$\geq\int_{0}^{T}\langle f_{0},\mu_{t}-P_{t}^{u_{0}}\rangle\,dt+\langle g_{0},\tilde{\mu}_{T}-\tilde{P}_{T}^{u_{0}}\rangle.$
Taking the infimum over $(f,g)\in\tilde{\mathcal{D}}$, we have
$V^{*}\geq\int_{0}^{T}\langle f_{0},\mu_{t}-P_{t}^{u_{0}}\rangle\,dt+\langle g_{0},\tilde{\mu}_{T}-\tilde{P}_{T}^{u_{0}}\rangle>-\infty.$
Hence $V^{*}$ is finite.
Step (ii). By (B1) and the bilinearity of the integral $\langle\cdot,\cdot\rangle$, clearly, the function
$\Psi[u,(f,g)]=\int_{0}^{T}\langle f,P_{t}^{u}-\mu_{t}\rangle\,dt+\langle g,\tilde{P}_{T}^{u}-\tilde{\mu}_{T}\rangle,$
is concavelike in $\tilde{\mathcal{D}}$, i.e., for any $(f_{1},g_{1}),(f_{2},g_{2})\in\tilde{\mathcal{D}}$, $\alpha\in[0,1]$, there exists $(f,g)\in\tilde{\mathcal{D}}$ such that $\alpha\Psi[u,(f_{1},g_{1})]+(1-\alpha)\Psi[u,(f_{2},g_{2})]\leq\Psi[u,(f,g)]$ for any $u\in\mathcal{U}$. Moreover, the function $\Psi[u,(f,g)]$ is convexlike in $\mathcal{U}$, i.e., for any $u,\tilde{u}\in\mathcal{U}$, $\alpha\in[0,1]$, there exists $u^{\prime}\in\mathcal{U}$ such that $\alpha\Psi[u,(f,g)]+(1-\alpha)\Psi[\tilde{u},(f,g)]\geq\Psi[u^{\prime},(f,g)]$ for any $(f,g)\in\tilde{\mathcal{D}}$. Indeed, we can always take an optimal $u_{f,g}$ for $u^{\prime}$. Also, by (B1), the set $\{\phi_{f,g}:(f,g)\in\tilde{\mathcal{D}}\}$ is compact. Hence, applying Sion’s minimax theorem [23, Theorem 4.1], we get
$\sup_{(f,g)\in\tilde{\mathcal{D}}}\inf_{u\in\mathcal{U}}\Psi[u,(f,g)]=\inf_{u\in\mathcal{U}}\sup_{(f,g)\in\tilde{\mathcal{D}}}\Psi[u,(f,g)],$
whence
$-V^{*}$ $=\sup_{\lambda>0}\sup_{(f,g)\in\tilde{\mathcal{D}}}\inf_{u\in\mathcal{U}}\left\{J(u;f_{0},g_{0})+\lambda\Psi[u,(f,g)]\right\}$
$=\sup_{\lambda>0}\inf_{u\in\mathcal{U}}\hat{J}_{\lambda}(u)$
where
$\hat{J}_{\lambda}(u)$ $=\int_{0}^{T}\langle f_{0}(t,\cdot,\cdot),P_{t}^{u}-\mu_{t}\rangle\,dt+\langle g_{0},P_{T}^{u}-\tilde{\mu}_{T}\rangle$
$\quad+\lambda\rho(P^{u},\mu),\quad u\in\mathcal{U}.$
For any $\varepsilon>0$ and $\lambda>0$,
$\inf_{u\in\mathcal{U}}\hat{J}_{\lambda}(u)\leq\inf_{u\in\mathcal{U}_{\rho,\varepsilon}}\Psi[u,(f_{0},g_{0})]+\lambda\varepsilon.$
Letting $\varepsilon\to 0$ and then taking the supremum with respect to $\lambda$ in the inequality just above, we get
$V^{*}\leq\lim_{\varepsilon\to 0}\inf_{u\in\mathcal{U}_{\rho,\varepsilon}}\Psi[u,(f_{0},g_{0})].$
Step (iii). Let $\varepsilon>0$ be fixed. Let $\{\lambda_{n}\}_{n=1}^{\infty}$ be a positive sequence such that $\lim_{n\to\infty}\lambda_{n}=0$ and
$\lim_{n\to\infty}\inf_{u\in\mathcal{U}}\hat{J}_{\lambda_{n}}(u)=\sup_{\lambda>0}\inf_{u\in\mathcal{U}}\hat{J}_{\lambda}(u).$
Let $\{\varepsilon_{n}\}_{n=1}^{\infty}$ be a positive sequence such that $\varepsilon_{n}\to 0$, $n\to\infty$. Then take $u_{n}\in\mathcal{U}$ satisfying
$\hat{J}_{\lambda_{n}}(u_{n})\leq\inf_{u\in\mathcal{U}}\hat{J}_{\lambda_{n}}(u)+\varepsilon_{n}$
for any $n\in\mathbb{N}$. We will show that
(III.2) $\lim_{n\to\infty}\lambda_{n}\rho(P^{u_{n}},\mu)=0.$
Assume contrary that
$\limsup_{n\to\infty}\lambda_{n}\rho(P^{u_{n}},\mu)=5\delta$
holds for some $\delta>0$. Then there exists a subsequence $\{n_{k}\}$ such that $\lim_{k\to\infty}\lambda_{n_{k}}\rho(P^{u_{n_{k}}},\mu)=5\delta$. Our assumption in the theorem means the existence of $u^{*}\in\mathcal{U}$ such that $\rho(P^{u^{*}},\mu)=0$. Thus $\hat{J}_{\lambda_{n}}(u_{n})\leq\hat{J}(u^{*})+\varepsilon_{n}$ for any $n$. In particular, the sequence $\{\hat{J}_{\lambda_{n_{k}}}(u_{n_{k}})\}_{k=1}^{\infty}$ is bounded, whence we can take a further subsequence $\{n_{k_{m}}\}$ such that
$\lim_{m\to\infty}\hat{J}_{\bar{\lambda}_{m}}(\bar{u}_{m})=\kappa:=\limsup_{k\to\infty}\hat{J}_{\lambda_{n_{k}}}(u_{n_{k}})<\infty,$
where $\bar{\lambda}_{m}=\lambda_{n_{k_{m}}}$ and $\bar{u}_{m}=u_{n_{k_{m}}}$. Then put $\bar{\gamma}_{m}=\gamma(P^{\bar{u}_{m}},\mu)$. Now choose $m_{0}$ and $m_{1}$ such that $\kappa<\hat{J}_{\bar{\lambda}_{m_{0}}}(\bar{u}_{m_{0}})+\delta$, $\hat{J}_{\bar{\lambda}_{m_{1}}}(\bar{u}_{m_{1}})<\kappa+\delta$, $\bar{\lambda}_{m_{1}}>7\bar{\lambda}_{m_{0}}$, and that $3\delta+\bar{\varepsilon}_{m_{0}}<\bar{\lambda}_{m_{1}}\bar{\gamma}_{m_{1}}<7\delta$, where $\bar{\varepsilon}_{m}=\varepsilon_{n_{k_{m}}}$. With these choices it follows that
$\kappa$ $<\hat{J}_{\bar{\lambda}_{m_{0}}}(\bar{u}_{m_{0}})+\delta\leq\inf_{u\in\mathcal{U}}\hat{J}_{\bar{\lambda}_{m_{0}}}(u)+\bar{\varepsilon}_{m_{0}}+\delta$
$\leq\hat{J}_{\bar{\lambda}_{m_{0}}}(\bar{u}_{m_{1}})+\bar{\varepsilon}_{m_{0}}+\delta$
$=\int_{0}^{T}\langle f_{0}(t,\cdot,\cdot),P_{t}^{\bar{u}_{m_{1}}}-\mu_{t}\rangle\,dt+\langle g_{0},\tilde{P}_{T}^{\bar{u}_{m_{1}}}-\tilde{\mu}_{T}\rangle$
$\quad+\frac{\bar{\lambda}_{m_{0}}}{\bar{\lambda}_{m_{1}}}\bar{\lambda}_{m_{1}}\bar{\gamma}_{m_{1}}+\bar{\varepsilon}_{m_{0}}+\delta.$
Observe
$\bar{\lambda}_{m_{0}}\bar{\gamma}_{m_{1}}+\bar{\varepsilon}_{m_{0}}+\delta<(1/7)\bar{\lambda}_{m_{1}}\bar{\gamma}_{m_{1}}+\bar{\varepsilon}_{m_{0}}+\delta$
$<2\delta+\bar{\varepsilon}_{m_{0}}<\bar{\lambda}_{m_{1}}\bar{\gamma}_{m_{1}}-\delta,$
whence $\kappa<\hat{J}_{\bar{\lambda}_{m_{1}}}(\bar{u}_{m_{1}})-\delta$, which is impossible.
In particular, $\rho(P^{u_{n}},\mu)\leq\varepsilon$ for a sufficiently large $n$. This yields
$\inf_{u\in\mathcal{U}}\hat{J}_{\lambda_{n}}(u)\geq\hat{J}_{\lambda}(u_{n})-\varepsilon_{n}$
$\geq\inf_{u\in\mathcal{U}_{\rho,\varepsilon}}\Psi[u,(f_{0},g_{0})]+\lambda_{n}\rho(P^{u_{n}},\mu)-\varepsilon_{n}.$
Letting $n\to\infty$ and then $\varepsilon\to 0$, we get
$V^{*}\geq\lim_{\varepsilon\searrow 0}\inf_{u\in\mathcal{U}_{\rho,\varepsilon}}\Psi[u,(f_{0},g_{0})].$
Therefore the theorem follows. ∎
Schrödinger problem
Here, we shall discuss a connection between Theorem III.2 and the Schrödinger bridge problems. Consider the case where the forward control problem (I.1) is described as
$J(u)=\mathbb{E}\int_{0}^{1}|u_{t}|^{2}dt,$
with $U=\mathbb{R}^{d}$, where the controlled process $X_{t}^{u}$ is governed by
$dX_{t}=u_{t}dt+dB_{t},\quad X_{0}\sim\tilde{\mu}_{0}.$
The class $\mathcal{D}$ of latent objectives is assumed to be given by
$\mathcal{D}=\{(|u|^{2},g):g\in C_{b}(\mathbb{R}^{d})\}.$
Consider $\mathcal{C}=\left\{g\in C_{b}(\mathbb{R}^{d}):\|g\|_{BL}\leq 1\right\}$, where $\|g\|_{BL}=\sup_{x\in\mathbb{R}^{d}}|g(x)|+\sup_{x\neq y}|g(x)-g(y)|/|x-y|$. By the well-known density arguments, we have
$\mathcal{D}=(|u|^{2},0)+\bigcup_{\lambda>0}\lambda\tilde{\mathcal{D}}$
with $\tilde{\mathcal{D}}=\{(0,g):g\in\mathcal{C}\}$. Thus (B1) is satisfied. Moreover,
$\rho(\pi,\nu)=\sup_{g\in\mathcal{C}}\,\langle g,\tilde{\pi}_{T}-\tilde{\nu}_{T}\rangle,$
whence $\rho$ defines a metric on $\mathcal{P}(\mathbb{R}^{d})$. More precisely, $\rho(\pi,\nu)=0$ if and only if $\tilde{\pi}_{T}=\tilde{\nu}_{T}$ (see, e.g., Dudley [6]). Consequently, our optimal transport problem becomes
(III.3) $\inf_{u\in\mathcal{U}_{\rho,0}}J(u)=\inf\left\{\mathbb{E}\int_{0}^{T}|u_{t}|^{2}dt:X_{T}^{u}\sim\tilde{\mu}_{T},\;u\in\mathcal{U}\right\},$
which is the so-called Schrödinger’s bridge problem.
Now, using Theorem III.2 we shall derive the duality formula for the Schrödinger’s problem, which is obtained in Mikami and Thieullen [15]. See Mikami [13], [14] for further generalizations. It should be emphasized that our approach is completely different from ones used in [15], [13], and [14]. Recall from Section II that we have assumed
$\int_{\mathbb{R}^{d}}|x|^{2}\tilde{\mu}_{0}(dx)<\infty.$
Theorem III.3 ([15]).
Suppose that $\tilde{\mu}_{T}$ has a positive density $\phi_{T}$ such that
$\int_{\mathbb{R}^{d}}(|y|^{2}+\log\phi_{T}(y))\phi_{T}(y)dy<\infty.$
Then we have
$\sup_{g\in C_{b}(\mathbb{R}^{d})}\left\{\inf_{u\in\mathcal{U}}\mathbb{E}\left[\int_{0}^{1}|u_{t}|^{2}dt+g(X_{1}^{u})\right]-\langle g,\tilde{\mu}_{1}\rangle\right\}$
$=\inf\left\{\mathbb{E}\int_{0}^{1}|u_{t}|^{2}dt:X_{1}^{u}\sim\tilde{\mu}_{1},\;u\in\mathcal{U}\right\}.$
Sketch of the proof.
Applying classical results in theory of the Schördinger problem, we can show that there exists $\{u_{t}^{*}\}_{0\leq t\leq T}\in\mathcal{U}$ such that $\inf_{u\in\mathcal{U}_{\rho,0}}J(u)=J(u^{*})$, whence $u^{*}$ is optimal to the Schrödinger’s problem (III.3). See, e.g. Nutz [20] and Jamison [9].
Let us show
(III.4) $\lim_{\varepsilon\searrow 0}\inf_{u\in\mathcal{U}_{\rho,\varepsilon}}J(u)=\inf_{u\in\mathcal{U}_{\rho,0}}J(u).$
Let $\varepsilon_{n}$ be a positive sequence such that $\varepsilon_{n}\searrow 0$, and $u^{(n)}\in\mathcal{U}_{\rho,\varepsilon_{n}}$ such that $J(u^{(n)})\leq\inf_{u\in\mathcal{U}_{\rho,\varepsilon_{n}}}J(u)+\varepsilon_{n}$. Put $X^{(n)}=X^{u^{(n)}}$. By a routine argument in probability theory we can show that the sequence $\mathbb{P}(X^{(n)})^{-1}$, $n\in\mathbb{N}$, is tight. So there exist a subsequence $\{n_{k}\}$ and $\hat{P}$ such that $\mathbb{P}(X^{(n_{k})})^{-1}$ weakly converges to $P^{*}$, a path measure of a unique solution of the Schördinger problem. This means that each subsequence $\{\mathbb{P}(X^{n_{k}})^{-1}\}$ contain a further subsequence $\{\mathbb{P}(X^{n_{k_{j}}})^{-1}\}$ converging weakly to $P^{*}$. Applying Theorem 2.6 in [4], we deduce that $\{\mathbb{P}(X^{n_{k}})^{-1}\}$ converges weakly to $P^{*}$ as $n\to\infty$. By the lower semi-continuity of the relative entropy,
$J(u^{*})=\lim_{n\to\infty}J(u^{(n)})\leq\lim_{n\to\infty}\inf_{u\in\mathcal{U}_{\rho,\varepsilon}}J(u).$
So (III.4) holds. This together with our theorem III.2 concludes. ∎
Acknowledgements
This study is supported by JSPS KAKENHI Grant Number JP17K05359.